A piano wire with mass 3.45 g and length 75.0 cm is stretched with a tension of 25.0 N . A wave with frequency 100 Hz and amplitude 1.50 mm travels along the wire. Calculate the average power carried by the wave. What happens to the average power if the wave amplitude is halved?
Answer #1

miu = mass per unit length
= 3.45*10^-3 Kg/0.75 m
= 4.6*10^-3 Kg/m

velocity,
v = sqrt (T/miu)
= sqrt (25/(4.6*10^-3))
= 73.72 m/s

Average power,
P = 0.5*miu*(2*pi*f)^2*A^2*v
= 0.5*(4.6*10^-3)*(2*pi*100)^2*(1.5*10^-3)^2*73.72
= 0.151 w
Answer: 0.151 w
--------------------------------------------------------------------
If amplitude is halfed, power will become 1/4
P = 0.151/4 = 0.038 w
Answer: 0.038 w